OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -2) and a directrix at y = 2

OpenStudy (imstuck):

If the focus is at (0,-2) and the directrix is the line y = 2, then the vertex is right between them, at (0,0). Because the directrix is above the focus the graph opens downward, so it is a y = x^2 parabola.

OpenStudy (anonymous):

mhmm

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

but

OpenStudy (imstuck):

So then, so far we have \[x ^{2}=4py\]because the origin is the vertex. Now we need to find p.

OpenStudy (imstuck):

p is the number of units the focus is away from the vertex. I graphed the points you provided and it appears that the focus is down 2 units from the vertex, so p = -2.

OpenStudy (imstuck):

Let me correct myself, quick-like...this opens downward, so the equation is this:\[x ^{2}=-4py\]i failed to put the - sign above.

OpenStudy (anonymous):

-2 still right :)

OpenStudy (imstuck):

So anyways, now we know that p = -2 (this is why the parabola opens downward, cuz of the negative), so filling in our equation with what we know...

OpenStudy (anonymous):

(x-h)^2=4p(y-k)

OpenStudy (imstuck):

\[x ^{2}=4(-2)y\]or\[x ^{2}=-8y\]

OpenStudy (imstuck):

That's it.

OpenStudy (imstuck):

TY for the medal.

OpenStudy (anonymous):

:) thats not an answer choice @IMStuck

OpenStudy (anonymous):

y2 = -2x y2 = -8x y equals negative 1 divided by 8 x squared y equals negative 1 divided by 2 x squared

OpenStudy (mathmate):

It would be equivalent to \(y=-\frac{x^2}{8} \)

OpenStudy (anonymous):

ahh because you isolate the y

OpenStudy (anonymous):

Wish I could give you a medal thanks you

OpenStudy (mathmate):

No, that's perfectly alright. You're most welcome! :)

OpenStudy (imstuck):

I medalled mathmate for you. You use my answer and solve it for y. If I would have known the form your answers were in, I could have done that part, too! Sorry!

OpenStudy (anonymous):

You can always do it from the definition \[ x^2 + (2 + y)^2 = (y - 2)^2\\ x^2 =-8 y\\ y=-\frac x 8 \] Since the parabola is the set of points equidistant from the focus and the directrix.

OpenStudy (anonymous):

How would I find the vertex of this qeuation?

OpenStudy (anonymous):

(0,0)

OpenStudy (anonymous):

To find it what steps would I use

OpenStudy (anonymous):

Is it i between the focus and directrix/?

OpenStudy (anonymous):

Yes, it is midpoint between the two

OpenStudy (anonymous):

Thank you that clarifies a lot,

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

So if I have a focus at (-8, 0) and a directrix at x = 8 then te midpoint is (0,0) and it's a horizontal parabola opening the the left correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you I fully understand it now. I found the answer to the next problem :D

OpenStudy (anonymous):

Great